OpenStudy (anonymous):

SystemEquations: solve x and y Solve x – 2y – 3z = 3 3x + y + z = 12 3x – 2y – 4z = 15

OpenStudy (anonymous):

Answers: (0, 0, -1) (3, 12, 15) (3, 9, -6) (-3, -9, 6)

OpenStudy (anonymous):

[1], [2], [3], etc. denote whether it is the first, second or third equation, etc. \[\large x – 2y – 3z = 3[1]\] \[\large 3x + y + z = 12[2]\] \[\large 3x – 2y – 4z = 15[3]\] \[[1]\times 3\] \[\large 3(x – 2y – 3z)=3(3)\] \[\large 3x-6y-9z=9[4]\] \[[2]-[3]\] \[\large (3x + y + z)-(3x – 2y – 4z)=(12)-(15)\] \[\large y+2y+z+4z=-3\] \[\large 3y+5z=-3[5]\] \[[2]-[4]\] \[\large (3x + y + z)-(3x-6y-9z)=(12)-(9)\] \[\large y+6y+z+9z=3\] \[\large 7y+10z=3[6]\] \[[5]\times 2\] \[\large 2(3y+5z)=2(-3)\] \[\large 6y+10z=-6[7]\] \[[6]-[7]\] \[\large (7y+10z)-(6y+10z)=(3)-(-6)\] \[\large 7y-6y=3+6\] \[\large y=9\] Sub the y-value into either equation [6] or [7] to find z. \[\large 7y+10z=3[6]\] \[\large 7(9)+10z=3\] \[\large 10z=3-63\] \[\large 10z=-60\] \[\large z=-6\] To find x, sub in the y and z values into either equation [1], [2], [3] or [4].

OpenStudy (anonymous):

gaussian elimination

OpenStudy (anonymous):

Thanks Man!

OpenStudy (anonymous):

But you don't need to find x to know which option is the correct one to choose from.

OpenStudy (anonymous):

Agreed

OpenStudy (anonymous):

do you know you understand how to get the answer ...?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

:)